Grade 128 grade maths

Using identities, evaluate.
8.9^2 ?????????????????????

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

8.9^2 = (9 - 0.1)^2
= 9^2- 1.8 + 0.1^2
= 81 - 1.8 + 0.01
= 79.21

Thanks